# The Relative Entropy of Expectation and Price

### v6

@@ V.1 Classical study
Three cases are considered in the studies with $n=$100$$ (left column), $1000$ (middle column) and $10000$ (right column), for the call option $f[x]=x^{+}$ (top two rows) and the digital option $f[x]=1_{x>0}$ (bottom two rows). The derivative security cannot be exactly replicated by the funding and hedging securities, and entropic risk optimisation compensates for this model risk with a bid-offer around the mid price and a skew to the hedge ratio. In the study with $n=$100$$ samples, the support is contained within the interval (-2.5,2.5), and there can be no option value outside this range of strikes. Increasing the number of samples widens the support and brings the solution closer to the limit of the standard normal variable.
@@
V.2 Quantum study
The quantum studies of figure 12 assume $q=0$ and $B$ and $Q$ are matrices with $n=2$ eigenvalues:
@@ V.2 Quantum study
The gap $\epsilon$ separates the eigenvalues of the funding security and the angle $\theta$ rotates the eigenbasis of the hedging security. The matrix $P=f[Q-kB]$ is obtained by first diagonalising the matrix $Q-kB$ and then applying the function $f$ to its eigenvalues:
@@ VI Entropic XVA
Figure 16: In this study, two counterparties have differing expectations for the future state. The first counterparty models the final underlying price as a uniform variable with zero mean and unit standard deviation. The second counterparty models the final underlying price as a normal variable with mean $\mu$ and unit standard deviation. The trade is then a call option on the final underlying price with zero strike. In each graph, the entropy-adjusted mean return on capital, computed with risk aversion $\alpha=1$, is plotted for both counterparties as a function of the option price per unit notional, including a range of values for the mean $\mu$ assumed by the second counterparty. The first counterparty is fully capitalised against default, and each row includes cases $c=\infty,0.5$ for the capital of the second counterparty. Each column includes cases $\lambda=0,1,2$ for the notional, where the case $\lambda=0$ corresponds to mid-pricing.
      The trade is viable when the entropy-adjusted mean return on capital is positive for both counterparties. The first counterparty buys the option when its price is low, and the second counterparty sells the option when its price is high. The second counterparty is more likely to sell when the mean $\mu$ is lower, as this reduces the expected option payoff. Model risk increases as the notional $\lambda$ is increased, which pushes the threshold prices in opposite directions and makes the trade less viable. Finally, for the second counterparty, decreasing the capital $c$ increases the risk of default. The CVA cost to the first counterparty and DVA benefit to the second counterparty then pushes both threshold prices lower.
VI.1 Margined derivatives
@@
As shown in figures 15, 15 and 15, net of all considerations the capital returned from the derivative is:
@@
00 0 $=\mathbb{E}_{1}^{\phi_{1}}[(dq_{1}-q_{1}r_{1}\,dt)\exp[-\alpha_{1}(dc_{1}-\delta_{1}\cdot dq_{1})]]$ (100)
00 0 $=\mathbb{E}_{2}^{\phi_{2}}[(dq_{2}-q_{2}r_{2}\,dt)\exp[-\alpha_{2}(dc_{2}-\delta_{2}\cdot dq_{2})]]$

### v7

@@ VII.1 Classical study
Three cases are considered in the studies with n=100n=\$100\$ (left column), $1000$ (middle column) and $10000$ (right column), for the call option $f[x]=x^{+}$ (top two rows) and the digital option $f[x]=1_{x>0}$ (bottom two rows). The derivative security cannot be exactly replicated by the funding and hedging securities, and entropic risk optimisation compensates for this model risk with a bid-offer around the mid price and a skew to the hedge ratio. In the study with n=100n=\$100\$ samples, the support is contained within the interval (-2.5,2.5), and there can be no option value outside this range of strikes. Increasing the number of samples widens the support and brings the solution closer to the limit of the standard normal variable.
@@
VII.2 Quantum study
The quantum studies of figure˜14 assume $q=0$ and $B$ and $Q$ are matrices with $n=2$ eigenvalues:
@@ VII.2 Quantum study
The gap $\epsilon$ separates the eigenvalues of the funding security and the angle $\theta$ rotates the eigenbasis of the hedging security. The matrix P=f[QkB]P=f[Q-kB] is obtained by first diagonalising the matrix QkBQ-kB and then applying the function $f$ to its eigenvalues:
@@ VIII Entropic XVA
Figure 18: In this study, two counterparties have differing expectations for the future state. The first counterparty models the final underlying price as a uniform variable with mean $0.5$ and unit standard deviation, and the second counterparty models the final underlying price as a normal variable with zero mean and unit standard deviation. The trade is then a call option on the final underlying price with zero strike. In each graph, the entropy-adjusted mean return on capital, computed with risk aversion $\alpha=1$, is plotted for both counterparties as a function of the option price per unit notional. The first counterparty is fully capitalised against default, and each row includes cases $c=\infty,0.5$ for the capital of the second counterparty. Each column includes cases $\lambda=0,1,2$ for the notional, where the case $\lambda=0$ corresponds to mid-pricing.
  The trade is viable when the entropy-adjusted mean return on capital is positive for both counterparties. The first counterparty is axed to buy the option when its price is low, and the second counterparty is axed to sell the option when its price is high. Model risk increases as the notional $\lambda$ is increased, which pushes the threshold prices in opposite directions and makes the trade less viable. Decreasing the capital $c$ increases the risk of default for the second counterparty. The CVA cost to the first counterparty and DVA benefit to the second counterparty then pushes both threshold prices lower.
VIII.1 Margined derivatives
@@
As shown in figures˜17, 17 and 17, net of all considerations the capital returned from the derivative is:
@@
$0$ $=\mathbb{E}_{1}^{\phi_{1}}[(dq_{1}-q_{1}r_{1}\,dt)\exp[-\alpha_{1}(dc_{1}-\delta_{1}\cdot dq_{1})]]$ (156)
$0$ $=\mathbb{E}_{2}^{\phi_{2}}[(dq_{2}-q_{2}r_{2}\,dt)\exp[-\alpha_{2}(dc_{2}-\delta_{2}\cdot dq_{2})]]$